# Ratchet current and scaling properties in a nontwist mapping

### v1

@@ I Introduction
About two decades ago, a new feature was observed in the transport of particles in the chaotic component of phase space in Hamiltonian systems: the existence of a preferential direction for the transport without an external bias, the so-called Hamiltonian ratchet [38, 39, 40, 41, 42, 43, 44]. The ratchet effect is defined by a directed current, or ratchet current, which is a preferential direction for the transport of particles in phase space without an external bias or directed force. This phenomenon has applications on a variaty for physical systems such as Josephson junctions [45, 46], Brownian [47] and moleculars [48] motors, cold atom systems [49, 50, 51, 52], and eletronic transport in superlattice [53, 54], to cite a few. Inital studies focused on the role of the external noise [55, 56, 47], which was latter replaced by a determistic chaotic dynamics with inertia terms in the equations of motion [57, 58, 59, 60]. Nonetheless, even purely Hamiltonian dynamics with mixed [38, 39, 41, 40, 43, 44] or completly chaotic [42] phase space can generate ratchet current. It was demonstrated that the ratchet effect is a consequence of spatial and/or temporal symmetry breaking of the system [39, 41, 44, 61]. Furthermore, Mugnaine et al. [62] demonstrated that due to a symmetry breaking in the extended standard nontwist mapping [63, 64], the twin island scenario no longer exists, impling a ratchet current in phase space due to the emergence of an unbalanced stickiness in different regions in phase space.
@@
In this paper, we study transport properties and diffusion for a different nontwist mapping, introduced in the context of $\mathbf{E}\times\mathbf{B}$ drift in toroidal plasmas [65, 66, 67], in the transition from integrability (zero perturbation) to non-integrability (small perturbation) when the phase space is limited. Firstly, we analyze the survival probability by introducing two exits symmetrically apart from the line $I=0$ and show that as long as there are no stability islands within the survival regions, the survival probability follows an exponential decay, and the decay rate scales as a power law with the limits of the survival regions. Furthermore, the survival probability exhibits scaling invariance with respect to the limits of the survival regions. Secondly, we consider two different ensembles of initial conditions, both with $\langle I\rangle\approx 0$ initially. The first ensemble is uniformly distributed along the whole survival region and we calculate the escape times and the escape basins for different survival region limits. We find that the measure of the bottom basin increases with the survival region area due to the inclusion of stability islands within this region. This ensemble, however, might be biased due to the distribution of initial conditions in phase space. To avoid this potential bias, we consider a second ensemble randomly distributed in a small region around the line $I=0$ and calculate the fraction of particles that escape from the top and bottom exits. We find, once again, that as we change the size of the survival region and islands start to appear, there is a tendency for the particles to escape through the bottom exit. Also, we observe a nonzero space average of the action, $\langle I\rangle$, characteristic of the ratchet effect.
@@
This paper is organized as follows. In Section LABEL:sec:mapping we describe the mapping under study and present some of its properties. In Section II we study the transport properties of the chaotic component of phase space. We begin by calculating the survival probability for different survival regions and in the sequence, we investigate whether the system exhibits unbiased transport by considering two different ensembles of initial conditions with zero initial average action. In Section III we present a phenomenological description of diffusion in the chaotic component of phase space. We obtain the critical exponents and compare our results with similar results in the literature. Section IV contains our final remarks.
@@
where $(I_{i},\theta_{i})$ is the canonical action-angle variables, $H_{0}$ is the integrable term and $\varepsilon H_{1}$ is a perturbation to the integrable system, with $\varepsilon$ controlling the transition from integrability ($\varepsilon=0$) to non-integrability ($\varepsilon>0$). The solution of Eq. (1) is a four-dimensional flow. However, since $\partialderivative*{H}{t}=0$, the Hamiltonian equals the total mechanical energy of the system, $H=E=T+V$, and it is a constant of motion. This allows us to eliminate one of the variables, e.g. $I_{2}$, from $H$, making it possible to write $H=H(I_{1},\theta_{1},\theta_{2},E)$. Therefore, orbits with energy $E$ are restricted to lie on a three-dimensional energy surface on the four-dimensional phase space. It is possible to decrease the dimension even further by considering an appropriate Poincaré section. We choose the plane $\theta_{1}\times I_{1}$ with $\theta_{2}$ constant as our Poincaré section, resulting in a two-dimensional mapping described by the following equations:
@@
Considering $p(\theta_{n},I_{n+1})\equiv 0$ and $h(\theta_{n},I_{n+1})=\sin(\theta_{n})$, and changing $K(\theta_{n})$ we obtain different systems well known in the literature, such as
@@
In this paper, we consider the following nontwist mapping introduced in the context of $\mathbf{E}\times\mathbf{B}$ drift in toroidal plasmas [65, 66, 67]
@@
$+\rho\frac{E(I_{n+1})}{\sqrt{\absolutevalue{I_{n+1}}}},$
where $q(I)$, $E(I)$, and $v(I)$ are given by, respectively,
@@
Although this mapping has many parameters, our only control parameter is the perturbation $\varepsilon$ and the remaining parameters are chosen accordingly to Refs. [66, 67] and can be found in Table 1. For the integrable case without pertubation ($\epsilon=0$), the variable $I$ is positive. However, with the pertubation $I$ can reach small negative values.
@@
Similarly to the Leonel mapping [74, 75, 76, 77, 78, 35], when the action $I_{n}$ is small, the angles $\theta_{n}$ and $\theta_{n+1}$ becomes uncorrelated due to the divergence of the second equation in Eq. (4), and thus producing chaotic regions for nonzero perturbation. For larger values of the action, the angles become correlated and regularity appears as stability islands and invariant spanning curves. In Fig. 1 we show the phase space of the mapping (4) for different perturbation values. For $\varepsilon=0$ (not shown) the mapping reduces to a nontwist radial map due to the nonmonotonicity of the functions $q(I)$, $E(I)$, and $v(I)$. In this case, the mapping is regular, the system is integrable, $I_{n}$ is constant, and the phase space has only periodic and quasiperiodic structures. On the other hand, for $\varepsilon>0$ the regularity is broken, even for $\varepsilon\ll 1$, and the phase space becomes mixed. There is a coexistence of chaotic and regular domains and as $\varepsilon$ increases, the chaotic region expands in the vertical direction due to the breaking of the KAM curves. For the chosen parameter values, the phase space is limited, i.e., and the transport in the vertical direction is confined to a finite area. Interestingly, the limits of the chaotic component shown in Fig. 1 are not symmetric with respect to the $I=0$ line.
@@
where $u_{i}$ is the eigenvector corresponding to the $i$th eigenvalue of $\mathbf{Df}^{n}$. A $d$-dimensional system has $d$ characteristic Lyapunov exponent, and we say the system is chaotic if at least one of them is positive. Furthermore, for Hamiltonian systems, which preserve volume in phase space under time evolution (Liouville’s theorem) [1], the sum of all Lyapunov exponents must equal to zero. Hence, for our two-dimensional area-preserving mapping, there are two characteristic exponents and they satisfy $\lambda_{+}\equiv\lambda_{1}=-\lambda_{2}$. In this case, all regular orbits, i.e., periodic, or quasi-periodic have zero Lyapunov exponents for infinite times, while chaotic orbits exhibit $\lambda_{+}>0$. In Fig. 2 we show the largest Lyapunov exponent, $\lambda_{+}$, as a function of the control parameter $\varepsilon$ for a single initial condition. We change $\varepsilon$ in a large interval and notice that $\lambda_{+}$, on the other hand, does not change significantly when compared to the range of variation of $\varepsilon$. This nearly constant value of $\lambda_{+}$ indicates that the chaotic component in phase space is scaling invariant with respect to the control parameter $\varepsilon$ [78]. We study such scaling properties in Sections II and III.
II Survival probability and ratchet current
In this section, we explore the transport of particles in the chaotic component of phase space. We have seen in the previous Section that the phase space for $\varepsilon\ll 1$ is limited, and therefore the particle is confined within a finite area. We consider two exits placed symmetrically from the $I=0$ line, and they define the survival region, $(\theta,I)\in[0,1)\times[-I_{\mathrm{esc}},I_{\mathrm{esc}}]$ (Fig. 3). We consider an ensemble of $M=10^{6}$ particles randomly distributed in $(\theta,I)\in[0,1)\times[-1.0\times 10^{-10},1.0\times 10^{-10}]$. We evolve each particle to at most $N=10^{6}$ times, and if the particle reaches one of the exits, i.e., if $I_{n}=\pm I_{\mathrm{esc}}$, it escapes and we interrupt the evolution of this particle and initialize another particle. We repeat this procedure until the whole ensemble is exhausted. From this ensemble, we calculate the survival probability, $P(n)$, that corresponds to the probability of a particle surviving along the dynamics in a given chaotic domain without escaping. In other words, it corresponds to the fraction of particles that have not yet reached one of the exits until the $n$th iteration. Mathematically, it is defined as $P(n)=N_{\mathrm{surv}}(n)/M$, where $N_{\mathrm{surv}}(n)$ is the number of particles that have survived until the $n$th iteration. The behavior of the survival probability depends strongly on the characteristics of phase space. For fully chaotic systems, the survival probability follows an exponential decay [29, 31, 32, 34] given by
@@ II Survival probability and ratchet current
Essentially, when some observable of a dynamical system exhibits scaling invariance, its expected behavior remains consistent and robust regardless of scale, i.e., we can rescale the system conveniently such that after a reparametrization, the observable is scale independent and exhibits universal features [84]. The scaling invariance of the survival probability, for instance, has been explored in a variaty of systems, such as the Leonel mapping [35], billiard systems [85], and, more recently, on fractional versions of the stardard map [86, 87].
Let us now investigate how transport occurs for particles uniformly distributed on the survival regions $(\theta,I)\in[0,1]\times[-I_{\mathrm{esc}},I_{\mathrm{esc}}]$ for $I_{\mathrm{esc}}=0.01,0.02,0.03,0.04$. Note that $\langle I\rangle\approx 0$ at $n=0$ for this ensemble. We consider $1080\times 1080$ initial conditions and iterate each one for at most $N=10^{6}$ times. We count the time each particle takes to reach either of the exits (top row of Fig. 5). Black and blue colors indicate a fast escape while the gray color corresponds to particles that have not yet escaped within $N=10^{6}$ iterations. The intermediate values of $T_{\mathrm{esc}}$ correspond to trapped particles. We observe that in cases of small survival regions, where stability islands are absent or nearly absent, the escape is fast. On the other hand, for the larger survival region considered [Fig. 5(d1)], most particles remain inside it for long times, with emphasis on the initial conditions near the stability islands (red color).
@@
However, how the particles are distributed in phase space might influence our previous conclusion. For example, some particles could be closer or farther away from the exits, while others might remain trapped for longer duration. To eliminate this potential bias, we consider an ensemble of $10^{6}$ randomly chosen initial conditions on the interval $(\theta,I)\in[0,1]\times[-1.0\times 10^{-10},1.0\times 10^{-10}]$, with $\langle I\rangle\approx 0$. Each particle is iterated up to $N=10^{6}$ times. We compute the fraction of particles that escape through either the bottom or top exit as a function of the survival region limits, $I_{\mathrm{esc}}$, for $\varepsilon=1.0\times 10^{-3}$, $\varepsilon=2.0\times 10^{-3}$, and $\varepsilon=3.0\times 10^{-3}$ (Fig. 6). For small survival regions, we observe that the escapes are evenly distributed between both the bottom and top exits. As $I_{\mathrm{esc}}$ increases, there is a tendency for the particles to escape through the bottom exit, more prominently for $\varepsilon=1.0\times 10^{-3}$ [Fig. 6(a)], characterizing the ratchet effect.
@@
Therefore, we have enough numerical evidence to state that the Hamiltonian nontwist mapping, given by Eq. (4), exhibits the Hamiltonian ratchet effect. To understand the cause of this directed transport, we consider a single chaotic initial condition, $(\theta,I)=(0.5,1.0\times 10^{-10})$, and iterate it for $N=10^{9}$ times with $\varepsilon=1.0\times 10^{-3}$, $\varepsilon=2.0\times 10^{-3}$, and $\varepsilon=3.0\times 10^{-3}$. We count the time it spends in both upper ($I>0$) and lower ($I<0$) regions of phase space. From this, we obtain two collections of recurrence times $\{t_{j}^{(\mathrm{upper})}\}_{j=1,2,\ldots,N_{t}^{(\mathrm{upper})}}$ and $\{t_{j}^{(\mathrm{lower})}\}_{j=1,2,\ldots,N_{t}^{(\mathrm{lower})}}$, where $N_{t}^{(\mathrm{upper})}$ and $N_{t}^{(\mathrm{lower})}$ are the number of recurrence times for the upper and lower regions, respectively  [^1] For the mentioned number of iterations, we obtained $N_{t}\sim 10^{6}$ for all values of $\varepsilon$., and define the probability distribution of recurrence times as $P(t)$ for both collections of recurrence times. Alternatively, we define the cumulative distribution of recurrence times, $Q(\tau)$, as follows:
$Q(\tau)=\sum_{t>\tau}P(t)=\frac{N_{\tau}}{N_{t}},$ (9)
where $N_{\tau}$ is the number of recurrence times larger than $\tau$, i.e., $t>\tau$. Figure 7(b) shows the cumulative distribution of recurrence times for the three values of $\varepsilon$ mentioned above for both collections of recurrence times, where the full lines correspond to the upper regions while the dashed lines to the lower one. We observe a power law tail for larger $\tau$, characteristic of systems that exhibit the stickiness effect [19, 20, 21, 22, 23, 24]. Furthermore, we observe different distributions for the upper and lower recurrence regions [full and dashed lines, respectively, in Fig. 7(b)]: the decay is faster for the upper region. The difference between these two distributions becomes more evident when we calculate the higher moments of $P(t)$:
@@
$\varepsilon$ $\langle t^{(\mathrm{U})}\rangle$ $\langle t^{(\mathrm{L})}\rangle$ $\langle t^{(\mathrm{L})}\rangle-\langle t^{(\mathrm{U})}\rangle$
@@
Therefore, the difference in the upper and lower distribution is evidence of an unbalanced stickiness in phase space. This phenomenon makes chaotic orbits to sticky in the upper and lower region of phase space unevenly due to the spatial asymmetry with respect to the $I=0$ line, creating the directed transport. Similar results have been reported in Ref. [62] for the extended standard nontwist map.
III Critical exponents and scaling law

### v2

@@ I Introduction
About two decades ago, a new feature was observed in the transport of particles in the chaotic component of phase space in Hamiltonian systems: the existence of a preferential direction for the transport without an external bias, the so-called Hamiltonian ratchet [38, 39, 40, 41, 42, 43, 44]. The ratchet effect is defined by a directed current, or ratchet current, which is a preferential direction for the transport of particles in phase space without an external bias. This phenomenon has applications on a variety for physical systems such as Josephson junctions [45, 46], Brownian [47] and moleculars [48] motors, cold atom systems [49, 50, 51, 52], and eletronic transport in superlattice [53, 54], to cite a few. Inital studies focused on the role of the external noise [55, 56, 47], which was later replaced by a determistic chaotic dynamics with inertia terms in the equations of motion [57, 58, 59, 60]. Nonetheless, even purely Hamiltonian dynamics with mixed [38, 39, 41, 40, 43, 44] or completly chaotic [42] phase space can generate ratchet current. It was demonstrated that the ratchet effect is a consequence of spatial and/or temporal symmetry breaking of the system [39, 41, 44, 61]. Furthermore, Mugnaine et al. [62] demonstrated that due to a symmetry breaking in the extended standard nontwist mapping [63, 64], the twin island scenario no longer exists, implyng a ratchet current in phase space due to the emergence of an unbalanced stickiness in different regions in phase space.
The anomalous transport of particles in magnetized plasmas is heavily impacted by the $\mathbf{E}\times\mathbf{B}$ drift motion, caused by fluctuations in the plasma [65]. In the case of passive particle transport, where the particles do not alter the electric field, the problem is described by a Hamiltonian system [66]. Furthermore, given the divergence-free nature of magnetic fields, the field lines can be described by a two-dimensional area-preserving mapping, with respect to the surface of section of the torus at a fixed toroidal angle [67]. In this paper, we study transport properties and diffusion for a nontwist mapping, introduced in the context of $\mathbf{E}\times\mathbf{B}$ drift in toroidal plasmas [68, 69, 70], in the transition from integrability (zero perturbation) to non-integrability (small perturbation) when the phase space is bounded. In the context of magnetic confined plasmas, as in Tokamak devices [71, 72], the use of non-monotonic profiles for physical quantities is related to nontwist effects like the existence of a shearless curve that acts as a robust barrier preventing the chaotic orbits from escaping the plasma [73]. Nontwist Hamiltonian systems are characterized by complex dynamics, including the formation of transport barriers and intricate island structures. Understanding these chaotic transport processes is essential for predicting and controlling plasma behavior.
@@
Initially, we analyze the survival probability by introducing two exits symmetrically apart from the line $I=0$ and show that as long as there are no stability islands within the survival regions, the survival probability follows an exponential decay, and the decay rate scales as a power law with the limits of the survival regions. Furthermore, the survival probability exhibits scaling invariance with respect to the limits of the survival regions. Then, we consider two different ensembles of initial conditions, both with $\langle I\rangle\approx 0$ initially. The first ensemble is uniformly distributed along the whole survival region and we calculate the escape times and the escape basins for different survival region limits. We find that the measure of the bottom basin increases with the survival region area due to the inclusion of stability islands within this region. This ensemble, however, might be biased due to the distribution of initial conditions in phase space. To avoid this potential bias, we consider a second ensemble randomly distributed in a small region around the line $I=0$ and calculate the fraction of particles that escape from the top and bottom exits. We find, once again, that as we change the size of the survival region and islands start to appear, there is a tendency for the particles to escape through the bottom exit. Also, we observe a nonzero space average of the action, $\langle I\rangle$, characteristic of the ratchet effect.
@@
This paper is organized as follows. In Section II we describe the mapping under study and present some of its properties. In Section III we study the transport properties of the chaotic component of phase space. We begin by calculating the survival probability for different survival regions and in the sequence, we investigate whether the system exhibits unbiased transport by considering two different ensembles of initial conditions with zero initial average action. In Section IV we present a phenomenological description of diffusion in the chaotic component of phase space. We obtain the critical exponents and compare our results with similar results in the literature. Section V contains our final remarks.
II The model and its properties
@@
where $(I_{i},\theta_{i})$ are the canonical action-angle variables, $H_{0}$ is the integrable term and $\varepsilon H_{1}$ is a perturbation to the integrable system, with $\varepsilon$ controlling the transition from integrability ($\varepsilon=0$) to non-integrability ($\varepsilon>0$). The solution of Eq. (1) is a four-dimensional flow. However, since $\partialderivative*{H}{t}=0$, the Hamiltonian equals the total mechanical energy of the system, $H=E=T+V$, and it is a constant of motion. This allows us to eliminate one of the variables, e.g. $I_{2}$, from $H$, making it possible to write $H=H(I_{1},\theta_{1},\theta_{2},E)$. Therefore, orbits with energy $E$ are restricted to lie on a three-dimensional energy surface on the four-dimensional phase space. It is possible to decrease the dimension even further by considering an appropriate Poincaré section. We choose the plane $\theta_{1}\times I_{1}$ with $\theta_{2}$ constant as our Poincaré section, resulting in a two-dimensional mapping described by the following equations:
@@
Considering $p(\theta_{n},I_{n+1})\equiv 0$ and $h(\theta_{n},I_{n+1})=\sin(\theta_{n})$, and changing $K(I_{n+1})$ we obtain different systems well known in the literature, such as
@@
In this paper, we consider the drift motion $\mathbf{E}\times\mathbf{B}$ in a toroidal plasma, where the electric field is generated by the combination of an external radial field $E$ and electrostatic potential $\tilde{\phi}$ generated by fluctuations in the plasma. The equations of the drift motion $\mathbf{E}\times\mathbf{B}$ are given by the following nontwist, area-preserving mapping [68, 69, 70]
@@
$+\rho\frac{E(I_{n+1})}{\sqrt{\absolutevalue{I_{n+1}}}}\mod 1,$
where $q(I)$, $E(I)$, and $v(I)$ are nonmonotonic functions, representing the safety factor, external applied radial electric field, and toroidal plasma velocity, respectively. The non-monotonicity of these functions makes the system violate the twist condition, which implies different dynamical properties [85, 79]. The nonmonotonic functions are given by,
@@
Although this mapping has many parameters, our only control parameter is $\varepsilon$ and the remaining parameters are chosen accordingly to Refs. [69, 70] and can be found in Table 1. The parameter $\varepsilon$ is proportional to the amplitude of the electrostatic instabilities that generate the drift motion. Thus, is the control parameter of integrability of the system. For the integrable case without perturbation ($\varepsilon=0$), the variable $I$ is positive. However, with the perturbation $I$ can reach small negative values.
@@
Similarly to the Leonel mapping [80, 81, 82, 83, 84, 35], when the action $I_{n}$ is small, the angles $\theta_{n}$ and $\theta_{n+1}$ becomes uncorrelated due to the divergence of the second equation in Eq. (4), and thus producing chaotic regions for nonzero perturbation. For larger values of the action, the angles become correlated and regularity appears as stability islands and invariant spanning curves. In Fig. 1 we show the phase space of the mapping (4) for different perturbation values. For $\varepsilon=0$ (not shown) the mapping reduces to a nontwist radial map due to the nonmonotonicity of the functions $q(I)$, $E(I)$, and $v(I)$. In this case, the mapping is regular, the system is integrable, $I_{n}$ is constant, and the phase space has only periodic and quasiperiodic structures. On the other hand, for $\varepsilon>0$ the regularity is broken, even for $\varepsilon\ll 1$, and the phase space becomes mixed. There is a coexistence of chaotic and regular domains and as $\varepsilon$ increases, the chaotic region expands in the vertical direction due to the breaking of the KAM curves. For the chosen parameter values, the phase space is bounded, i.e., and the transport in the vertical direction is confined to a finite area. Interestingly, the limits of the chaotic component shown in Fig. 1 are not symmetric with respect to the $I=0$ line.
@@
where $u_{i}$ is the eigenvector corresponding to the $i$th eigenvalue of $\mathbf{Df}^{n}$. A $d$-dimensional system has $d$ characteristic Lyapunov exponent, and we say the system is chaotic if at least one of them is positive. Furthermore, for Hamiltonian systems, which preserve volume in phase space under time evolution (Liouville’s theorem) [1], the sum of all Lyapunov exponents must equal zero. Hence, for our two-dimensional area-preserving mapping, there are two characteristic exponents and they satisfy $\lambda_{+}\equiv\lambda_{1}=-\lambda_{2}$. In this case, all regular orbits, i.e., periodic, or quasi-periodic have zero Lyapunov exponents for infinite times, while chaotic orbits exhibit $\lambda_{+}>0$. In Fig. 2 we show the largest Lyapunov exponent, $\lambda_{+}$, as a function of the control parameter $\varepsilon$ for a single initial condition. We change $\varepsilon$ in a large interval and notice that $\lambda_{+}$, on the other hand, does not change significantly when compared to the range of variation of $\varepsilon$. This nearly constant value of $\lambda_{+}$ indicates that the chaotic component in phase space is scaling invariant with respect to the control parameter $\varepsilon$ [84]. We study such scaling properties in Sections III and IV.
III Survival probability and ratchet current
In this section, we explore the transport of particles in the chaotic component of phase space. We have seen in the previous Section that the phase space for $\varepsilon\ll 1$ is bounded, and therefore the particle is confined within a finite area. We consider two exits placed symmetrically from the $I=0$ line, and they define the survival region, $(\theta,I)\in[0,1)\times[-I_{\mathrm{esc}},I_{\mathrm{esc}}]$ (Fig. 3). We consider an ensemble of $M=10^{6}$ particles randomly distributed in $(\theta,I)\in[0,1)\times[-1.0\times 10^{-10},1.0\times 10^{-10}]$. We evolve each particle to at most $N=10^{6}$ times, and if the particle reaches one of the exits, i.e., if $I_{n}=\pm I_{\mathrm{esc}}$, it escapes and we interrupt the evolution of this particle and initialize another particle. We repeat this procedure until the whole ensemble is exhausted. From this ensemble, we calculate the survival probability, $P(n)$, that corresponds to the probability of a particle surviving along the dynamics in a given chaotic domain without escaping. In other words, it corresponds to the fraction of particles that have not yet reached one of the exits until the $n$th iteration. Mathematically, it is defined as $P(n)=N_{\mathrm{surv}}(n)/M$, where $N_{\mathrm{surv}}(n)$ is the number of particles that have survived until the $n$th iteration. The behavior of the survival probability depends strongly on the characteristics of phase space. For fully chaotic systems, the survival probability follows an exponential decay [29, 31, 32, 34] given by
@@ III Survival probability and ratchet current
Essentially, when some observable of a dynamical system exhibits scaling invariance, its expected behavior remains consistent and robust regardless of scale, i.e., we can rescale the system conveniently such that after a reparametrization, the observable is scale independent and exhibits universal features [91]. The scaling invariance of the survival probability, for instance, has been explored in a variety of systems, such as the Leonel mapping [35], billiard systems [92], and, more recently, on fractional versions of the standard map [93, 94].
Let us now investigate how transport occurs for particles uniformly distributed on the survival regions $(\theta,I)\in[0,1)\times[-I_{\mathrm{esc}},I_{\mathrm{esc}}]$ for $I_{\mathrm{esc}}=0.01,0.02,0.03,0.04$. Note that $\langle I\rangle\approx 0$ at $n=0$ for this ensemble. We consider $1080\times 1080$ initial conditions and iterate each one for at most $N=10^{6}$ times. We count the time each particle takes to reach either of the exits (top row of Fig. 5). Black and blue colors indicate a fast escape while the gray color corresponds to particles that have not yet escaped within $N=10^{6}$ iterations. The intermediate values of $T_{\mathrm{esc}}$ correspond to trapped particles. We observe that in cases of small survival regions, where stability islands are absent or nearly absent, the escape is fast. On the other hand, for the larger survival region considered [Fig. 5(d1)], most particles remain inside it for long times, with emphasis on the initial conditions near the stability islands (red color).
@@
However, how the particles are distributed in phase space might influence our previous conclusion. For example, some particles could be closer or farther away from the exits, while others might remain trapped for a longer duration. To eliminate this potential bias, we consider an ensemble of $10^{6}$ randomly chosen initial conditions on the interval $(\theta,I)\in[0,1]\times[-1.0\times 10^{-10},1.0\times 10^{-10}]$, with $\langle I\rangle\approx 0$. Each particle is iterated up to $N=10^{6}$ times. We compute the fraction of particles that escape through either the bottom or top exit as a function of the survival region limits, $I_{\mathrm{esc}}$, for $\varepsilon=1.0\times 10^{-3}$, $\varepsilon=2.0\times 10^{-3}$, and $\varepsilon=3.0\times 10^{-3}$ (Fig. 6). For small survival regions, we observe that the escapes are evenly distributed between both the bottom and top exits. As $I_{\mathrm{esc}}$ increases, there is a tendency for the particles to escape through the bottom exit, more prominently for $\varepsilon=1.0\times 10^{-3}$ [Fig. 6(a)], characterizing the ratchet effect.
@@
Therefore, we have enough numerical evidence to state that the Hamiltonian nontwist mapping, given by Eq. (4), exhibits the Hamiltonian ratchet effect. To understand the cause of this directed transport, we consider a single chaotic initial condition, $(\theta,I)=(0.5,1.0\times 10^{-10})$, and iterate it for $N=10^{9}$ times with $\varepsilon=1.0\times 10^{-3}$, $\varepsilon=2.0\times 10^{-3}$, and $\varepsilon=3.0\times 10^{-3}$. We count the time it spends in both upper ($I>0$) and lower ($I<0$) regions of phase space. From this, we obtain two collections of recurrence times $\{t_{j}^{(\mathrm{U})}\}_{j=1,2,\ldots,N_{t}^{(\mathrm{U})}}$ and $\{t_{j}^{(\mathrm{L})}\}_{j=1,2,\ldots,N_{t}^{(\mathrm{L})}}$, where $N_{t}^{(\mathrm{U})}$ and $N_{t}^{(\mathrm{L})}$ are the number of recurrence times for the upper and lower regions, respectively  [^1] For the mentioned number of iterations, we obtained $N_{t}\sim 10^{6}$ for all values of $\varepsilon$., and define the probability distribution of recurrence times as $P^{(\textrm{U,L})}(t)$ for both collections of recurrence times. Alternatively, we define the cumulative distribution of recurrence times, $Q^{(\textrm{U,L})}(\tau)$, as follows:
$Q^{(\textrm{U,L})}(\tau)=\sum_{t>\tau}P^{(\textrm{U,L})}(t)=\frac{N_{\tau}^{(\textrm{U,L})}}{N_{t}^{(\textrm{U,L})}},$ (9)
where $N^{(\textrm{U,L})}_{\tau}$ is the number of recurrence times larger than $\tau$, i.e., $t>\tau$ for the respective regions. Figure 7(b) shows the cumulative distribution of recurrence times for the three values of $\varepsilon$ mentioned above for both collections of recurrence times, where the full lines correspond to the upper regions while the dashed lines to the lower one. We observe a power law tail for larger $\tau$, characteristic of systems that exhibit the stickiness effect [19, 20, 21, 22, 23, 24]. Furthermore, we observe different distributions for the upper and lower recurrence regions [full and dashed lines, respectively, in Fig. 7(b)]: the decay is faster for the upper region. The difference between these two distributions becomes more evident when we calculate the higher moments of $P(t)$:
@@
$\varepsilon$ $\langle t\rangle^{(\mathrm{U})}$ $\langle t\rangle^{(\mathrm{L})}$ $\langle t\rangle^{(\mathrm{L})}-\langle t\rangle^{(\mathrm{U})}$
@@
Therefore, the difference in the upper and lower distribution is evidence of an unbalanced stickiness in phase space. This phenomenon makes chaotic orbits to sticky in the upper and lower region of phase space unevenly due to the spatial asymmetry with respect to the $I=0$ line, creating the directed transport. Similar results have been reported in Ref. [62] for the extended standard nontwist map. It is important to emphasize, however, that is the existence of asymmetric islands that leads to such a phenomenon, i.e., there is no symmetry transformation between the upper and lower islands. Additionally, the perturbation $\varepsilon$ and the survival region also influence the transport (Fig. 6). In regions without stability islands, the particles escape evenly from the bottom and top exits. Thus, the inherent asymmetry by itself is not enough to generate the ratchet current in our model.
IV Critical exponents and scaling law